OpenStudy (swaqqout_kid1):

Which value for x makes the sentence true? 3x = -3x + 24 x = 8 x = 3 x = 0 x = 4 @ttb123456789

OpenStudy (swaqqout_kid1):

@SolomonZelman @Secret-Ninja

OpenStudy (swaqqout_kid1):

@magepker728

OpenStudy (secret-ninja):

All you have to do is switch out x for the numbers. Lets try the first one together.

OpenStudy (secret-ninja):

3 x 8 = -3 x 8 +24 Do you know how to start?

OpenStudy (swaqqout_kid1):

subtrat 3x from 3

OpenStudy (secret-ninja):

You have to use PEMDAS (Parenthesis, exponents, multiplication, division, addition, subtraction) and you have to multiply first, in this case.

OpenStudy (secret-ninja):

3 x 8 is 24 -3 x 8 = -24 24 = -24 +24

OpenStudy (secret-ninja):

Now you have to add. -24 +24 = 0 24=0 is what you get, but 24 is not 0, so A is not it.

OpenStudy (secret-ninja):

Do you think you can do it or do you want me to stay?

OpenStudy (secret-ninja):

@SwaqqOut_KID1 If you want help, you have to stay and see what I show you, otherwise, you wont be able to do it, okay?

OpenStudy (swaqqout_kid1):

k

OpenStudy (secret-ninja):

So do you think you got this, or do you want more help for the other ones?

OpenStudy (swaqqout_kid1):

B

OpenStudy (texaschic101):

he's guessing

OpenStudy (swaqqout_kid1):

no it waz a typo i meant 2 put D a$$hole

OpenStudy (secret-ninja):

Lets test it out 3 x 3 = -3 x 3 + 24 You multiply (using PEMDAS) 9 = -81 + 24 You add now. 9 = -57. So B is wrong.

OpenStudy (texaschic101):

excuse me...no reason to call me that. It is D

OpenStudy (secret-ninja):

Do not use foul language, otherwise, I will report you and not help you at all.

OpenStudy (texaschic101):

very good explaining secret ninja

OpenStudy (secret-ninja):

Lets try D out, if you are so insistant. 3x = -3x + 24 3 x 4 = -3 x 4 + 24 12 = -12 + 24 12 = 12 God job, D is correct.

OpenStudy (secret-ninja):

Thank you @texaschic101

OpenStudy (secret-ninja):

Do you think you can remember this process? It can help much further on in school. @SwaqqOut_KID1

OpenStudy (secret-ninja):

P - Parenthesis ( ) E - Exponents \[0^{0}\] M - Multiplication x D - Division / A - Addition + S - Subtraction -

OpenStudy (secret-ninja):

Make sure to close the question before you go. :) Thanks. \(\Large\cal\color{darkorchid}{Have~a~nice~day!}\)

OpenStudy (anonymous):

x=4